For what value of k does the following system have infinitely many solutions?
kx + y + z=0
x + 2y + kz = 0
-X+3z=0

Found 2 solutions by greenestamps, ikleyn:
Answer by greenestamps(13200)



Solve the 3rd equation for z and substitute in the other two equations.

x = 3z

3kz+y+z=0
3z+2y+kz=0

y+(3k+1)z=0
2y+(3+k)z=0

There are infinitely many solutions if the two equations are equivalent.

2y+(6k+2)z=0
2y+(3+k)z=0

6k+2=3+k
5k=1
k=1/5

ANSWER: k=1/5


Answer by ikleyn(52785)

.

This system has infinitely many solutions if and only if the determinant of the
coefficient matrix is zero.

So, calculate the determinant of the matrix as a function of "k" and equate it to zero.

From this equation, find the value of "k".